Question 18 (1 point) Determine the probability. The deck of cards in CandyLand has 48 single-move cards, 12 double-move cards and 6 character cards. If a card is drawn at random, what is the probability it is a double-move? 12 1/3
Answer #1

here as total number of cards =48+12+6=66

number of double move cards =12

hence P(double move card)=12/66 =2/11